OpenStudy (anonymous):

Use separation of variables to solve the initial value problem. dy/dx = x/y and y =2 when x =1. So far I've multiplied both sides by dx and got dy/y = x/y * y dx then did the integral stuff and got ∫1/y dy = ∫x dx which equals ∫y^-1 dy = ∫x dx and that equals -1/y^2 = 1 am I doing this right so far? and if i am, how do i continue this?

OpenStudy (kainui):

Nope. You need to multiply both sides by "dx" and both sides by "y" to get everything separated out. ydy=xdx That should be what you get from the algebra, right?

OpenStudy (anonymous):

Oh okay. Was looking at my teachers example from this lesson and must have followed it wrong. so after that I should plug the numbers and and get 2 = 1 + c and that'd make C = 1, correct?

OpenStudy (kainui):

Well from this you're doing an integral... right? \[\frac{ 1 }{ 2 }y^2=\frac{ 1 }{ 2 }x^2 +C\] Plugging in we get \[\frac{ 1 }{ 2 }2^2=\frac{ 1 }{ 2 }1^2 +C\] \[2=\frac{ 1 }{ 2 } +C\] \[\frac{ 3 }{ 2 } =C\]

OpenStudy (kainui):

So now you can just rewrite it as: \[y^2=x^2+3\]

OpenStudy (anonymous):

Ah, alright. Thank you. I got a final answer of \[\sqrt{x^{2}+3}\] and that appears to be correct according to my book. Thanks again!

OpenStudy (kainui):

Yeah, and remember when you take a square root you can have both positive and negative. So if your initial conditions were: y=-2 and x=1 then you would do everything exactly the same except have: \[y=-\sqrt{x^2+3}\]